Bakery A sells bread for $2 per loaf that costs $0.80 per loaf to make. Bakery A gives a 75% discount for its bread at the end of the day. Demand for the bread is normally distributed with a mean of 300 and a standard deviation of 30. What order quantity maximizes expected profit for Bakery A?

a) 324.00

b) 300.17

c) 325.25
Answer #1

Salvage value = $2 × (1 - 75%) = $0.50.

Overage cost = $0.80 - $0.50 = $0.30.

Underage cost = $2 - $0.80 = $1.20.

The critical ratio = 1.2/(1.2 + 0.3) = 0.8.

z = Normsinv(0.8) = 0.84.

Order quantity = 300 + (0.84 × 30) = 325.25.

Hence the correct option is c. 325.25